# Interior Spectral Windows and Transport for Discrete Fractional Laplacians on $d$-Dimensional Hypercubic Lattices

### v1

@@ 3. Proof of main result
Now, let $\chi\in C_{c}^{\infty}(\mathcal{J})$. Since $W(Q)$ is compact on $\ell^{2}(\mathbb{Z}^{d})$ when $W(n)\to 0$, the Helffer-Sj-strand functional calculus yields $\chi(H)-\chi(H_{0})\in\mathcal{K}(\ell^{2}({\mathbb{Z}}^{d}))$. Therefore
@@ 4.1. Stationary representation of the scattering matrix on interior windows
as bounded maps $\mathcal{H}_{s}\to\mathcal{H}_{-s}$; here $R_{0}(z)=(H_{0}-z)^{-1}$.
The $T$-operator. Let $R(z)=(H-z)^{-1}$. Define
@@
as bounded maps $\mathcal{H}_{s}\to\mathcal{H}_{-s}$ for $\Im z\neq 0$, with a.e. limits $T(\lambda\pm\mathrm{i}0)$ on $\mathcal{I}$ by the localized LAP.
Wave matrices and the on-shell scattering matrix. Let $\chi\in C_{c}^{\infty}(\mathcal{I})$ equal $1$ on $\mathcal{I}_{0}\Subset\mathcal{I}$. The local wave operators
@@
(4.3) $W_{\pm}(\lambda)=\mathrm{id}-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda\pm\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*},\qquad\text{a.e. }\lambda\in\mathcal{I},$
@@
(4.4) $S(\lambda)=\mathrm{id}-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda+\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*},\qquad\text{a.e. }\lambda\in\mathcal{I}.$
Proof of (4.3) and (4.4).
By the localized LAP on $\mathcal{I}$ and Kato smoothness ([AmBoGe, Thms. 7.4.1-7.5.1]), the time-dependent representation
$\big{\langle}(\mathcal{W}_{\pm}-\mathrm{id})f,\,g\big{\rangle}=\mp\mathrm{i}\!\int_{0}^{\pm\infty}\big{\langle}e^{-\mathrm{i}tH}\,\chi(H_{0})\,W\,e^{\mathrm{i}tH_{0}}f,\,g\big{\rangle}\,\mathrm{d}t$
holds as a Bochner integral $\mathcal{H}_{s}\to\mathcal{H}_{-s}$ for $s>\tfrac{1}{2}$. Insert the spectral resolution of $H_{0}$ on $\mathcal{I}$ via $\mathcal{U}_{0}$ and use Sokhotski-Plemelj (justified by the LAP) to convert time integrals into boundary values of resolvents. Using (4.2) (and the optical identity (4.5)), one obtains
$\mathcal{U}_{0}(\mathcal{W}_{\pm}-\mathrm{id})\mathcal{U}_{0}^{*}:\ \phi(\lambda)\mapsto-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda\pm\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*}\,\phi(\lambda),$
which is (4.3); then (4.4) follows. ∎
Optical theorem, unitarity, and continuity. Using $R_{0}(z)-R_{0}(\bar{z})^{*}=(z-\bar{z})R_{0}(\bar{z})^{*}R_{0}(z)$ and (4.2) yields, for a.e. $\lambda\in\mathcal{I}$,
@@
as a form identity on $\mathcal{H}_{-s}$, $s>\tfrac{1}{2}$. Combining (4.1)-(4.5) gives $W_{\pm}(\lambda)^{*}W_{\pm}(\lambda)=\mathrm{id}$ and hence $S(\lambda)$ is unitary a.e. on $\mathcal{I}$. Moreover, by the localized LAP, $\lambda\mapsto S(\lambda)$ is strongly continuous on $\mathcal{I}$.
@@
With $\Gamma(\lambda+\mathrm{i}0):=(\mathrm{id}-R_{0}(\lambda+\mathrm{i}0)W)^{-1},$ we have $T(\lambda+\mathrm{i}0)=W\,\Gamma(\lambda+\mathrm{i}0)$ and
@@
Note that $\Gamma(\lambda+\mathrm{i}0)$ is not the full resolvent $R(\lambda+\mathrm{i}0)$.
@@ 4.2.2. Ballistic transport in time average
Fix nondecreasing $\Phi\in C_{c}^{\infty}(\mathbb{R})$ with $\Phi^{\prime}\geq 0$ and $\Phi^{\prime}\not\equiv 0$; set $F_{R}:=\Phi(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/R)$ and $f(t):=e^{-\mathrm{i}tH}\chi(H)f$. The commutator expansion ([AmBoGe, Prop. 6.2.10]) together with $H\in\mathcal{C}^{1,1}_{\mathrm{loc}}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}})$ on $\mathcal{I}$ yields, on $\mathrm{Ran}\,\chi(H)$,
$\mathrm{i}[H,F_{R}]=\frac{1}{R}\,\Phi^{\prime}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/R)\,\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]+\mathcal{O}(R^{-2})$
in $\mathcal{B}(\ell^{2}({\mathbb{Z}}^{d}))$, uniformly on $\mathcal{I}$ (localized calculus). Hence
@@
$\frac{d}{dt}\langle f(t),F_{R}f(t)\rangle=\frac{1}{R}\langle f(t),\Phi^{\prime}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/R)\,\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]\,f(t)\rangle+\mathcal{O}(R^{-2})\|f\|^{2}.$
Insert $\chi(H)$ on both sides of $\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]$ and use the localized Mourre estimate on $\mathcal{I}$ with compact remainder ([AmBoGe, Thm. 7.2.9]): $\chi(H)\,\mathrm{i}[H,]\,\chi(H)\geq c_{\mathcal{I}}\,\chi(H)^{2}-\chi(H)K_{\mathcal{I}}\chi(H).$ Integrating from 0 to $\mathcal{T}$ and using $\Phi^{\prime}\geq 0$ gives
$\frac{c_{\mathcal{I}}}{R}\!\int_{0}^{\mathcal{T}}\!\!\langle f(t),\Phi^{\prime}(A_{\mathbb{Z}^{d}}/R)f(t)\rangle dt\ \leq\ C\|f\|^{2}+\frac{1}{R}\!\int_{0}^{\mathcal{T}}\!\!\big{|}\langle f(t),\Phi^{\prime}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/R)K_{\mathcal{I}}f(t)\rangle\big{|}dt+CR^{-2}\|f\|^{2}.$
By Kato smoothness (from the LAP, [AmBoGe, Thm. 7.5.1]), the compact term is $\leq C\|f\|^{2}$, uniformly in $\mathcal{T},R$. Choosing $R=\alpha\mathcal{T}$ and $\Phi^{\prime}$ with $\Phi^{\prime}\geq c\,\mathbf{1}_{\{|x|\leq v_{\mathcal{I}}\}}$ yields
$\frac{1}{\mathcal{T}}\int_{0}^{\mathcal{T}}\|\mathbf{1}_{\{|A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}|\leq v_{\mathcal{I}}t\}}\,u(t)\|^{2}\,dt\ \leq\ \frac{C_{\mathcal{I}}}{\log(1+\mathcal{T})}\,\|u\|^{2},$
after the standard logarithmic optimization ([AmBoGe, $\S$7.4]). ∎
@@ 4.2.3. Continuity in the fractional exponents
The norm-resolvent continuity in $\vec{\mathbf{r}}$ implies inner–outer continuity of the spectra. This is in the spirit of results on magnetic families on $\mathbb{Z}^{d}$ [PR16] and of results for magnetic pseudodifferential families in the continuum [AMP10]. In particular, a Mourre estimate uniform on $\mathcal{I}$ yields a uniform LAP and the strong continuity of the scattering matrix $\lambda\mapsto S_{\vec{\mathbf{r}}}(\lambda)$ on $\mathcal{I}$.
@@
If $W\in\ell^{1}(\langle n\rangle^{1+\epsilon})$ for some $\epsilon>0$, then $\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}$ is finite (counting multiplicities).
@@
Let $\chi\in C_{c}^{\infty}(\mathcal{I})$. As in the proof of Theorem 4.2, $\chi(H)-\chi(H_{0})\in\mathfrak{S}_{1}$. By Helffer–Sj
"ostrand and functional calculus stability, $E_{\mathcal{I}}(H)-E_{\mathcal{I}}(H_{0})\in\mathfrak{S}_{1}$. Since $H_{0}$ is purely absolutely continuous on $\mathcal{I}$ (interior window), $E_{\mathcal{I}}(H_{0})$ has no point spectrum. Therefore $E_{\mathcal{I}}(H)$ has finite rank and $\big{(}\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}\big{)}<\infty$ (counting multiplicities). ∎
@@
The conclusion that $\bigl{(}\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}\bigr{)}<\infty$ follows from a uniform Mourre estimate on $\mathcal{I}$ and the limiting absorption principle (LAP), which yield compactness of the localized resolvent and thus only finitely many eigenvalues in $\mathcal{I}$ (each of finite multiplicity). In the continuum magnetic setting, Athmouni-Purice prove Schatten–von Neumann criteria within the magnetic Weyl calculus, allowing one to place the Birman-Schwinger type operators in $\mathfrak{S}_{p}$ and derive analogous spectral finiteness statements; see [AP18]. For general background on trace ideals/Birman–Schwinger and on the Mourre framework, see [Sim05],[ReSi, vol.IV].

### v2

@@ 3. Proof of main result
Now, let $\chi\in C_{c}^{\infty}(\mathcal{J})$. Since $W(Q)$ is compact on $\ell^{2}(\mathbb{Z}^{d})$ when $W(n)\to 0$, the Helffer–Sjstrand functional calculus yields $\chi(H)-\chi(H_{0})\in\mathcal{K}(\ell^{2}({\mathbb{Z}}^{d}))$. Therefore
@@ 4.1. Stationary representation of the scattering matrix on interior windows
as bounded maps $\mathcal{H}_{s}\to\mathcal{H}_{-s}$; here $R_{0}(z)=(H_{0}-z)^{-1}$. The $T$-operator. Let $R(z)=(H-z)^{-1}$. Define
@@
as bounded maps $\mathcal{H}_{s}\to\mathcal{H}_{-s}$ for $\Im z\neq 0$, with a.e. limits $T(\lambda\pm\mathrm{i}0)$ on $\mathcal{I}$ by the localized LAP. Wave matrices and the on-shell scattering matrix. Let $\chi\in C_{c}^{\infty}(\mathcal{I})$ equal $1$ on $\mathcal{I}_{0}\Subset\mathcal{I}$. The local wave operators
@@
(4.3) $W_{\pm}(\lambda)=\mathrm{id}_{\mathscr{H}_{\lambda}}-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda\pm\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*},\qquad\text{a.e. }\lambda\in\mathcal{I},$
@@
(4.4) $S(\lambda)=\mathrm{id}_{\mathscr{H}_{\lambda}}-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda+\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*},\qquad\text{a.e. }\lambda\in\mathcal{I}.$
Optical theorem, unitarity, and continuity.
Assume the localized LAP holds on $\mathcal{I}$, i.e.
$\sup_{\lambda\in\mathcal{I}}\,\big{\|}\langle A\rangle^{-s}R_{0}(\lambda\pm\mathrm{i}0)\langle A\rangle^{-s}\big{\|}+\sup_{\lambda\in\mathcal{I}}\,\big{\|}\langle A\rangle^{-s}R(\lambda\pm\mathrm{i}0)\langle A\rangle^{-s}\big{\|}<\infty,\quad s>\tfrac{1}{2},$
and that $T(z)$ is defined by (4.2) and admits boundary values $T(\lambda\pm\mathrm{i}0)\in\mathcal{B}(\mathcal{H}_{s},\mathcal{H}_{-s})$ for a.e. $\lambda\in\mathcal{I}$. Let $\Gamma_{0}(\lambda):\mathcal{H}_{s}\to\mathscr{H}_{\lambda}$ be the boundary trace operator (the “free Fourier transform on the energy shell”) satisfying Stone’s formula
(4.5) $\frac{1}{2\pi\mathrm{i}}\big{(}R_{0}(\lambda+\mathrm{i}0)-R_{0}(\lambda-\mathrm{i}0)\big{)}=\Gamma_{0}(\lambda)^{*}\Gamma_{0}(\lambda)\quad\text{in }\mathcal{B}(\mathcal{H}_{s},\mathcal{H}_{-s}),\quad s>\tfrac{1}{2}.$
Using the resolvent identity
$R_{0}(z)-R_{0}(\bar{z})=(z-\bar{z})\,R_{0}(\bar{z})R_{0}(z),$
and (4.2), one obtains for a.e. $\lambda\in\mathcal{I}$ the optical theorem
@@
as a form identity on $\mathcal{H}_{-s}$ ($s>\tfrac{1}{2}$).
Define the stationary wave operators on the shell by
$W_{\pm}(\lambda):=\mathrm{id}_{\mathscr{H}_{\lambda}}-2\pi\mathrm{i}\,\Gamma_{0}(\lambda)\,T(\lambda\pm\mathrm{i}0)\,\Gamma_{0}(\lambda)^{*}\quad\in\mathcal{B}\big{(}\mathscr{H}_{\lambda}\big{)},$
and the scattering matrix $S(\lambda):=W_{+}(\lambda)^{*}W_{-}(\lambda)$. Combining (4.5)–(4.6) yields for a.e. $\lambda\in\mathcal{I}$
$W_{\pm}(\lambda)^{*}W_{\pm}(\lambda)=\mathrm{id}_{\mathscr{H}_{\lambda}},\qquad S(\lambda)^{*}S(\lambda)=S(\lambda)S(\lambda)^{*}=\mathrm{id}_{\mathscr{H}_{\lambda}},$
so $S(\lambda)$ is unitary almost everywhere on $\mathcal{I}$.
Finally, by the localized LAP and the continuity of $\Gamma_{0}(\lambda)$ in $\lambda$ as a map $\mathcal{H}_{s}\to\mathscr{H}_{\lambda}$, the boundary values $T(\lambda\pm\mathrm{i}0)$ depend continuously on $\lambda$ in $\mathcal{B}(\mathcal{H}_{s},\mathcal{H}_{-s})$. Hence $\lambda\mapsto S(\lambda)$ is strongly continuous on $\mathcal{I}$.
@@
It is convenient to factor $T(z)$ as
$T(z)=W\,\Gamma(z),\qquad\Gamma(z):=(\mathrm{id}-R_{0}(z)W)^{-1}.$
In particular,
@@
We stress that $\Gamma(\lambda+\mathrm{i}0)$ is not the full interacting resolvent $R(\lambda+\mathrm{i}0)=(H-\lambda-\mathrm{i}0)^{-1}$; rather, it is an auxiliary operator that encodes the resummation of the Born series in the definition of $T(z)$.
@@ 4.2.2. Ballistic transport in time average
Set $u:=\chi(H)f$ and $f(t):=e^{-\mathrm{i}tH}u$. Pick $\Phi\in C^{\infty}(\mathbb{R})$, bounded and constant at both ends, such that $\Phi^{\prime}\in C_{c}^{\infty}(\mathbb{R})$, $\Phi^{\prime}\geq 0$, and $\Phi^{\prime}\not\equiv 0$. For $v>0$ (to be chosen later) and $t\geq 1$, define the time–dependent observable
$F_{t}:=\Phi\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}.$
Then, for $g(t):=\langle f(t),F_{t}f(t)\rangle$, one has
$\frac{d}{dt}g(t)=\big{\langle}f(t),\partial_{t}F_{t}f(t)\big{\rangle}\ +\ \mathrm{i}\,\big{\langle}f(t),[H,F_{t}]f(t)\big{\rangle}.$
The explicit time derivative reads
$\partial_{t}F_{t}\ =\ -\frac{1}{t}\,\frac{A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}}{vt}\,\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)},$
and, since $\Phi^{\prime}\geq 0$ and $\Phi$ is nondecreasing, one has $\langle f(t),\partial_{t}F_{t}f(t)\rangle\leq 0$. Thus, this term can be dropped in an upper bound.
@@
By the commutator expansion (see [AmBoGe, Prop. 6.2.10]) and the assumption $H\in\mathcal{C}^{1,1}_{\mathrm{loc}}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}})$ on $\mathcal{I}$, one has on $\mathrm{Ran}\,\chi(H)$:
$\mathrm{i}[H,F_{t}]\ =\ \frac{1}{vt}\,\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}\,\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]\ +\ \mathcal{O}(t^{-2})\quad\text{in }\mathcal{B}(\ell^{2}(\mathbb{Z}^{d})).$
Inserting $\chi(H)$ around $\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]$ and applying the localized Mourre estimate with compact remainder ([AmBoGe, Thm. 7.2.9]), there exist $c_{\mathcal{I}}>0$ and a compact operator $K_{\mathcal{I}}$ such that
$\chi(H)\,\mathrm{i}[H,A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}]\,\chi(H)\ \geq\ c_{\mathcal{I}}\,\chi(H)^{2}\ -\ \chi(H)K_{\mathcal{I}}\chi(H).$
Thus, for $t\geq 1$,
$\frac{d}{dt}g(t)$ $\ \geq\ \frac{c_{\mathcal{I}}}{vt}\,\big{\langle}f(t),\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}f(t)\big{\rangle}$
@@
$\quad-\ \frac{1}{vt}\,\big{|}\big{\langle}f(t),\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}K_{\mathcal{I}}f(t)\big{\rangle}\big{|}\ -\ C\,t^{-2}\,\|u\|^{2},$
with $C<\infty$ independent of $t$.
Integrating from $1$ to $\mathcal{T}$, the left–hand side telescopes and is bounded by $|g(\mathcal{T})-g(1)|\leq 2\|\Phi\|_{\infty}\|u\|^{2}$. Hence
(4.7) $\frac{c_{\mathcal{I}}}{v}\int_{1}^{\mathcal{T}}\frac{1}{t}\,\big{\langle}f(t),\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}f(t)\big{\rangle}\,dt\ \leq\ C_{1}\|u\|^{2}\ +\ \frac{1}{v}\int_{1}^{\mathcal{T}}\frac{1}{t}\,R(t)\,dt,$
with $R(t):=\big{|}\langle f(t),\Phi^{\prime}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt))K_{\mathcal{I}}f(t)\rangle\big{|}+C^{\prime}t^{-1}\|u\|^{2}$.
The compact term is handled via Kato smoothness stemming from the LAP ([AmBoGe, Thm. 7.5.1]). Decompose $K_{\mathcal{I}}=K_{\mathrm{fr}}+K_{\mathrm{sm}}$ (finite rank + small norm). By Cauchy–Schwarz and HH–smoothness on $\mathrm{Ran}\chi(H)$, one obtains
$\int_{1}^{\mathcal{T}}\frac{1}{t}\,\big{|}\langle f(t),\Phi^{\prime}(A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt))K_{\mathcal{I}}f(t)\rangle\big{|}\,dt\ \leq\ C_{2}\,\log(1+\mathcal{T})\,\|u\|^{2},$
uniformly in $v>0$; also $\int_{1}^{\mathcal{T}}t^{-2}\,dt\leq 1$. Inserted into (4.7), this yields
$\int_{1}^{\mathcal{T}}\frac{1}{t}\,\big{\langle}f(t),\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}f(t)\big{\rangle}\,dt\ \leq\ C_{3}\,\log(1+\mathcal{T})\,\|u\|^{2}.$
Finally, choose $\Phi$ such that $\Phi^{\prime}\geq c\,\mathbf{1}_{\{|x|\leq 1\}}$ for some $c>0$. Then
$\big{\langle}f(t),\Phi^{\prime}\!\big{(}A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}/(vt)\big{)}f(t)\big{\rangle}\ \geq\ c\,\big{\|}\mathbf{1}_{\{|A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}|\leq vt\}}\,f(t)\big{\|}^{2}.$
Therefore,
$\int_{1}^{\mathcal{T}}\frac{1}{t}\,\big{\|}\mathbf{1}_{\{|A_{\mathbb{Z}^{d},\vec{\mathrm{r}}}|\leq vt\}}\,e^{-\mathrm{i}tH}\chi(H)f\big{\|}^{2}\,dt\ \leq\ C_{4}\,\log(1+\mathcal{T})\,\|f\|^{2}.$
Since $\int_{1}^{\mathcal{T}}t^{-1}\,dt=\log\mathcal{T}$, the standard logarithmic optimization argument ([AmBoGe, §7.4]) gives, after division by $\mathcal{T}$ and with $v=v_{\mathcal{I}}$ chosen small enough, the desired inequality on $0,\mathcal{T}]$ (the interval $[0,1]$ being controlled trivially by $C\,\mathcal{T}^{-1}\|f\|^{2}$). This completes the proof. ∎
Remark 4.4 (On the choice of $\Phi$).
There is no nontrivial $\Phi\in C_{c}^{\infty}$ that is monotone. Instead, one chooses $\Phi\in C^{\infty}$ bounded, constant at both ends, with $\Phi^{\prime}\in C_{c}^{\infty}$, $\Phi^{\prime}\geq 0$. For instance, take $\psi\in C_{c}^{\infty}$, $\psi\geq 0$, and set $\Phi(x):=\int_{-\infty}^{x}\psi(s)\,ds$. After normalization, this satisfies $\Phi^{\prime}\geq c\,\mathbf{1}_{[-1,1]}$ and is perfectly suited for the argument.
@@ 4.2.3. Continuity in the fractional exponents
The norm-resolvent continuity in $\vec{\mathrm{r}}$ implies inner–outer continuity of the spectra. This is in the spirit of results on magnetic families on $\mathbb{Z}^{d}$ [PR16] and of results for magnetic pseudodifferential families in the continuum [AMP10]. In particular, a Mourre estimate uniform on $\mathcal{I}$ yields a uniform LAP and the strong continuity of the scattering matrix $\lambda\mapsto S_{\vec{\mathrm{r}}}(\lambda)$ on $\mathcal{I}$.
@@
If $W\in\ell^{1}(\langle n\rangle^{1+\epsilon})$ for some $\epsilon>0$, then $\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}$ is finite, counting multiplicities.
@@
Let $\chi\in C_{c}^{\infty}(\mathcal{I})$. As in the proof of Theorem 4.2, one has
$\chi(H)-\chi(H_{0})\in\mathfrak{S}_{1},$
where $\mathfrak{S}_{1}$ denotes the trace class. By the Helffer–Sjöstrand formula and stability of the functional calculus under trace class perturbations, it follows that
$E_{\mathcal{I}}(H)-E_{\mathcal{I}}(H_{0})\in\mathfrak{S}_{1}.$
Since $H_{0}$ is purely absolutely continuous on $\mathcal{I}$ (being inside the interior of the spectrum), the spectral projection $E_{\mathcal{I}}(H_{0})$ has no pure point component. Consequently, the compact perturbation above implies that $E_{\mathcal{I}}(H)$ differs from a projection of infinite rank by a trace class operator, hence $E_{\mathcal{I}}(H)$ has at most finite rank.
Now recall that the rank of $E_{\mathcal{I}}(H)$ is exactly the sum of the multiplicities of eigenvalues of $H$ in $\mathcal{I}$, i.e.
$\mathrm{rank}\,E_{\mathcal{I}}(H)\ =\ \sum_{\lambda\in\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}}m(\lambda),$
where $m(\lambda)$ is the algebraic multiplicity of $\lambda$. Since $E_{\mathcal{I}}(H)$ has finite rank, this sum is finite, which proves that
$\big{(}\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}\big{)}<\infty,$
counting multiplicities. ∎
@@
The conclusion that $\bigl{(}\sigma_{\mathrm{pp}}(H)\cap\mathcal{I}\bigr{)}<\infty$ follows from a uniform Mourre estimate on $\mathcal{I}$ and the limiting absorption principle (LAP), which yield compactness of the localized resolvent and thus only finitely many eigenvalues in $\mathcal{I}$ (each of finite multiplicity). In the continuum magnetic setting, Athmouni-Purice prove Schatten–von Neumann criteria within the magnetic Weyl calculus, allowing one to place the Birman–Schwinger type operators in $\mathfrak{S}_{p}$ and derive analogous spectral finiteness statements; see [AP18]. For general background on trace ideals/Birman–Schwinger and on the Mourre framework, see [Sim05],[ReSi, vol.IV].